Using the enthalpy changes given below, calculate the enthalpy change for the following equation:

N2(g) + 2O2(g) -----> 2 NO2

using the following equations:

2NO -----> N2 + O2 delta H + -180.0 kJ

NO2 -----> NO + 1/2 O2 delta H= 112.0 kJ
Answer #1

The given reactions are
2NO \rightarrow N2 + O2 \: \: \: \delta H = -180.0 kJ \\ NO2 \rightarrow NO + 1/2 O2 \: \: \: \delta H= 112.0 kJ

Reverse above reactions
N2 + O2 \rightarrow 2NO \: \: \: \delta H = 180.0 kJ......(1) \\ NO + 1/2 O2 \rightarrow NO2 \: \: \: \delta H= -112.0 kJ......(2)
Multiply equation (2) with 2
2NO + O2 \rightarrow 2NO2 \: \: \: \delta H= -112.0 \times 2 = - 224.0 kJ......(3)
Add equations (1) and (3)
N2 +2O2 \rightarrow 2NO2 \delta H = 180.0 - 224.0 = -44 kJ